# SPAGS: Sparse-View Articulated Object Reconstruction from Single State via Planar Gaussian Splatting

### v2

@@ 4.3 Ablation Study
Table 6: Ablation of the number of input views on PartNet-Mobility dataset. Mean results are reported.
Input Views CD $\downarrow$ CD-d $\downarrow$ CD-s $\downarrow$
@@
Ablation of The Number of Input Views. To verify the performance of our method for different numbers of input views, we conduct ablation experiments on the PartNet-Mobility dataset. Note that all input views are perceived by the optimal view perception method. The experimental results in Table. 6 prove that our method can enhance the performance of mesh reconstruction by using more input views. Moreover, we observe that when the input views are less than 4, the reconstruction quality suffers a sharp decline. This is because the information provided by too few views is insufficient, and high-fidelity mesh reconstruction through the insufficient image supervision poses a significant challenge.
Table 7: Ablation of key components on PartNet-Mobility dataset. Mean results are reported.
Settings CD $\downarrow$ CD-d $\downarrow$ CD-s $\downarrow$
@@
Ablation of Key Components. We conduct ablation of proposed core components on PartNet-Mobility dataset. In the “w/o planar Gaussians” setting, we employ the depth rendering pipeline from GaussianObject. In “w/o pseudo labels” setting, we omit the pseudo labels generated from reliable regions. The experimental results in Table. 7 demonstrate that each component makes a significant improvement to both whole surface and part-level reconstruction performance, especially for the planar Gaussians and the refinement. This confirms the effectiveness of proposed key components, and the superiority of planar Gaussians as well as the refinement via few-shot diffusion.
4.4 Training Time Analysis
@@
4.5 Generalization to the Real World
We present the quantitative results of real-world mesh reconstruction in Table. 4 and the corresponding qualitative results in Fig. 7. Our method significantly outperforms existing approaches in mean results across all metrics and demonstrates strong generalization on real-world data. As shown in Fig. 7, our method enables high-fidelity reconstruction for unseen real-world articulated objects with two or more parts.
Figure 7: The qualitative results of our real-world performance. Static parts are marked in blue, dynamic parts in yellow, and additional dynamic parts in red and green. Red arrows represent the joints.
4.6 Conclusion and Limitation
@@
The limitations of our method lie in the challenge of reconstructing articulated objects with transparent materials or extremely small size in real-world environments. Future work will focus on integrating physically based rendering to model transparent materials [transparentgs], and incorporating super-resolution techniques [xie2024supergs] to improve reconstruction with low-resolution image inputs.

### v4

@@ IV-C Ablation Study
TABLE VII: Ablation of the number of input views on PARIS dataset. Mean results are reported.
Input Views CD $\downarrow$ CD-m $\downarrow$ CD-s $\downarrow$
@@
Ablation of The Number of Input Views. To verify the performance of our method for different numbers of input views, we conduct ablation experiments on the PARIS dataset. Note that all input views are perceived by the optimal view perception method. The experimental results in Table. VII prove that our method can enhance the performance of mesh reconstruction by using more input views. Moreover, we observe that when the input views are less than 4, the reconstruction quality suffers a sharp decline. This is because the information provided by too few views is insufficient, and high-fidelity mesh reconstruction through the insufficient image supervision poses a significant challenge.
TABLE VIII: Quantitative results for mesh reconstruction on real-world data.
Metrics Method Stapler USB Scissor Knife Drawer Bin Mean
CD $\downarrow$ CoherentGS [26] 30.65 10.92 13.12 8.76 104.02 64.84 38.72
Sparse2DGS [34] 34.07 7.63 15.24 3.21 84.15 65.38 34.95
GaussianObject [40] 11.24 3.09 5.23 4.76 21.54 15.30 10.19
SPAGS (Ours) 3.13 2.34 5.76 2.82 16.36 4.00 5.74
F1 $\uparrow$ CoherentGS [26] 0.11 0.14 0.02 0.16 0.01 0.02 0.08
Sparse2DGS [34] 0.14 0.20 0.02 0.21 0.03 0.03 0.11
GaussianObject [40] 0.16 0.25 0.05 0.22 0.01 0.07 0.13
SPAGS (Ours) 0.18 0.24 0.10 0.27 0.03 0.07 0.15
EMD $\downarrow$ CoherentGS [26] 0.25 0.23 0.35 0.15 0.64 0.59 0.37
Sparse2DGS [34] 0.16 0.18 0.23 0.18 0.58 0.50 0.31
GaussianObject [40] 0.09 0.11 0.16 0.14 0.23 0.27 0.17
SPAGS (Ours) 0.06 0.07 0.10 0.07 0.13 0.08 0.09
Ablation of Key Components. We conduct ablation of proposed core components on PARIS dataset. In the “w/o planar Gaussians” setting, we employ the depth rendering pipeline from GaussianObject. In “w/o pseudo labels” setting, we omit the pseudo labels generated from reliable regions. The experimental results in Table. IX demonstrate that each component makes a significant improvement to both whole surface and part-level reconstruction performance, especially for the planar Gaussians and the refinement. This confirms the effectiveness of proposed key components, and the superiority of planar Gaussians as well as the refinement via few-shot diffusion.
TABLE IX: Ablation of key components on PARIS dataset. Mean results are reported.
Settings CD $\downarrow$ CD-m $\downarrow$ CD-s $\downarrow$
@@
IV-D Training Time Analysis
@@
Figure 12: Qualitative results of our method for real-world articulated objects.
IV-E Generalization to the Real World
To further evaluate our generalization to real-world objects, we reconstruct objects by capturing sparse images of indoor articulated objects under a static state. Since methods like ArtGS and REArtGS++ are inapplicable using single-state observations, we conduct quantitative comparisons for whole meshes against SOTA sparse reconstruction baselines in Tab. VIII, and provide qualitative results of articulated modeling in Fig. 12.
As shown in Tab. VIII, our method significantly outperforms existing approaches in mean results across all metrics and demonstrates strong generalization on real-world data. As shown in Fig. 12, our method enables high-fidelity reconstruction for unseen real-world articulated objects with two or more parts.
IV-F Conclusion and Limitation
@@
Despite its effectiveness, our method exhibits limitations when handling real-world articulated objects with transparent materials or minute geometric structures. Standard Gaussian representations struggle to capture complex optical phenomena like refraction, while extremely small parts suffer from insufficient pixel-level constraints in sparse settings. Moving forward, we intend to integrate physically based rendering (PBR) [6] into our pipeline to explicitly model transparent and specular surfaces. Furthermore, we aim to incorporate neural super-resolution techniques [37] to recover high-frequency details from low-resolution inputs, enabling robust reconstruction of micro-articulated structures.
Acknowledgements
This work was supported by National Natural Science Foundation of China under Grant 62302143.